In a sample of 1000 workers, if 200 of them are graduates, then what is the sample proportion of graduate workers? In addition, if 650 of them are married, what is the proportion of married workers?
1
Expert's answer
2020-05-18T00:39:00-0400

Sample proportion is given by "\frac{x}{n}"

Sample proportion for graduate workers is "\frac{200}{1000}=\frac{1}{5}"

Proportion of married workers is "\frac{650}{1000}=\frac{13}{20}"